OpenStudy (princeevee):

Calculate the discriminant to determine the number of real roots. y = x2 + 3x + 9 How many real roots does the equation have? one real root two real roots no real roots no solution

OpenStudy (welshfella):

For 2 real roots b^2 - 4ac > 0 for one real root b^2 - 4ac = 0 for b^2 - 4ac < 0 no real roots

OpenStudy (princeevee):

so c

OpenStudy (welshfella):

yes the discriminant = 9 - 36 = -27 so no real roots

OpenStudy (welshfella):

good job